# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 7.5.2 Gauge symmetry enlargement
As we have already discussed several times, see, e.g., Sec. 7.5.1, at transitions of LGW and LGW× type, gauge models may display an enlargement of the global symmetry. Analogously, at topological transitions, one may observe an enlargement of the gauge symmetry—this occurs in ${\mathbb{Z}}_{Q}$ gauge models, see Sec. 3.5.2. The models we are now going to discuss have instead been investigated to understand whether there are transitions with gauge symmetry enlargement, that are effectively described by a continuum GFT. In the present context, one is looking for models with discrete gauge symmetry that undergo charged transitions that are associated with the AHFT, with an enlarged U(1) gauge symmetry.
Charged transitions controlled by the AHFT occur in the $N$-component LAH model with noncompact Abelian gauge field (along the CH line), in the higher-charge ($Q>1$) $N$-component LAH models with compact Abelian gauge field (along the DC-OD line). In both cases the gauge group can be approximated with arbitrary accuracy by considering discrete compact or noncompact subgroups and one can define models that are only invariant under these discrete subgroups. The question is whether these discrete models also have charged transitions as their continuous counterparts.
Let us start by discussing the discretized version [369] of the noncompact $N$-component LAH model with Hamiltonian (45) defined in Sec. 5.1. The discretized model is defined in the same way, the only difference being the set of values that the gauge field $A_{{\bm{x}},\mu}$ takes. In the discretized model, the gauge field takes only the values $2\pi m/q$, with $m\in\mathbb{Z}$. The model is invariant under a restricted set of gauge transformations (we indicate with ${\mathbb{Z}}_{q}^{(nc)}$ the corresponding group)
@@
The original noncompact U(1) gauge theory is formally recovered in the limit $q\to\infty$. Note that the gauge group is discrete, but infinite. As already discussed in Sec. 5.1, a proper definition of the model requires the introduction of C boundary conditions.
Figure 18: Sketch of the phase diagram of the 3D $N$-component LAH model with discrete gauge symmetry. In the noncompact case the transition in $A$ is of the XY universality class, while the transitions at the points $B$ and $C$ are of the IXY universality class [369]. In the compact case the transition in $A$ is of the XY universality class, while the transitions at the points $B$ and $C$ are of the $\mathbb{Z}_{Q}$ and $\mathbb{Z}_{q}$ universality class, respectively [212]. Both in the compact and in the noncompact case we have $J_{c}(A),\kappa_{c}(C)\to\infty$ in the limit $q\to\infty$.
@@
Due to its larger global symmetry, the phase diagram of the discrete $N$-component LAH model is significantly more complex than that of the $N$-component LAH with U(1) gauge symmetry. A sketch is shown in Fig. 18. Using the same arguments presented in Sec. 6.2, it is possible to show (see Ref. [369] for more details) that an IXY transition occurs both for $J=\infty$ (as in the noncompact U(1) gauge AH model) and for $J=0$. Moreover, the transition on the $J=0$ axis occurs at $\kappa_{c}(J=0)=q^{2}\kappa_{c}(J=\infty)$, where $\kappa_{c}(J=\infty)=0.076051(2)$, so $\kappa_{c}(J=0)\to\infty$ for $q\to\infty$. This should not come as a surprise, since we should recover the phase diagram of the model with U(1) symmetry with no transition on the $J=0$ line, see Fig. 2, for $q\to\infty$. Moreover, two transitions are present on the $\kappa=0$ axis: a transition for $J=J_{\mathrm{SU}(N)}$, where the global SU($N$) symmetry is spontaneously broken, and a transition at $J=J_{q}$, associated with the spontaneous breaking of the U(1)$/\mathbb{Z}_{q}$ symmetry. As discussed above, we expect this symmetry to break in a phase in which the SU($N$) symmetry is already broken, implying $J_{q}>J_{\mathrm{SU}(N)}$ (therefore the point $A$ in Fig. 18 corresponds to $J=J_{q}$). Note that $J_{q}\to\infty$ as $q\to\infty$, since this transition is not present in the phase diagram of the noncompact U(1) gauge model, see Fig. 2.
In Ref. [369] the phase diagram of the $N$-component $\mathbb{Z}_{q}^{(nc)}$ gauge model has been numerically investigated, studying in detail the critical behavior along the transition line $M_{2}M_{3}$ in Fig. 18. Indeed, for $q\to\infty$, this line corresponds to the CH line (see Fig. 2) of the noncompact U(1) gauge model, which is the line where the AHFT critical behavior emerges if the number $N$ of scalar components is large enough, as discussed in Sec. 5.3.3. The numerical results show that, for $N=25$ and $q\geq 5$, the transitions along the line $M_{2}M_{3}$ are continuous, and that their critical behavior is the same as along the CH line in the noncompact U(1) gauge model. Thus, for $q\geq 5$ and large values of $N$, the transitions along the line $M_{2}M_{3}$ are effectively associated with the stable charged FP of the AHFT, with a corresponding enlargement of the gauge symmetry. On the right of the meeting point $M_{3}$ we instead have O($2N$)× critical transitions.
@@
In Ref. [212] the universality class of the transitions along the line $M_{2}M_{3}$ has been investigated for $Q=2$ and $N=15$. In the U(1) invariant model, recovered for $q\to\infty$, these transitions are charged transitions associated with the stable FP of the AHFT, see Sec. 6.3.2. On the other hand, only discontinuous transitions were identified along the $M_{2}M_{3}$ line for for values of $q$ up to $q=10$. Apparently, in the discretized model there are no charged transitions. It is still an open issue to understand why the two apparently similar discretized models, the compact and the noncompact model, behave differently. The origin may lie in the fact that in the noncompact model the gauge group is discrete but still infinite, while in the compact case the gauge group is discrete and finite.
@@ 8 Non-Abelian Higgs field theories
Non-Abelian gauge symmetries play a fundamental role in the construction of quantum and statistical field theories that describe phenomena in various physical contexts. In high-energy physics they provide the framework to formulate the fundamental theories of the strong and electroweak interactions, see, e.g., Refs. [1, 2, 370], while in condensed-matter physics they provide an effective description of classical and quantum critical phenomena characterized by the emergence of non-Abelian gauge fields, see, e.g., Refs [371, 96, 20, 372, 16]. In the presence of low-energy modes associated with scalar fields, they give rise to the non-Abelian Higgs mechanism [103, 104, 105], which, for instance, is one of the main features of the theory of electroweak interactions [1, 2, 370]. The emerging Higgs phases in non-Abelian Higgs (NAH) theories crucially depend on the interplay between local and global symmetries, and on the spontaneous breaking patterns of the global symmetry, see, e.g., Refs. [373, 258, 41, 374, 273, 62, 64, 65, 33, 375, 96, 97, 259, 376, 98, 99].
Three-dimensional non-Abelian gauge models can be used to effectively describe the critical behaviors of 3D statistical non-Abelian gauge models at finite-temperature phase transitions [377, 378, 379, 380, 381, 382, 383, 384, 385], and investigate some aspects of fundamental mechanisms, such as confinement [373, 386, 387, 256]. The peculiar critical behaviors developed by 3D NAH models also provide further possible scenarios for finite-temperature continuous transitions in $D=3+1$ non-Abelian gauge QFTs, where the temperature is related to the (inverse) finite size along the fourth Euclidean direction of their path-integral formulation, see, e.g., Refs. [25, 377, 378, 379, 380, 381, 382, 388, 383, 223, 224, 168, 26, 23, 169]. NAH theories have been also considered as effective theories of phenomena involving emerging non-Abelian gauge fields interacting with scalar modes, see, e.g., Refs. [371, 96], providing effective theories to describe some aspects of the phenomenology of high-temperature cuprate superconductors, see, e.g., Refs. [389, 390, 391].
In this section we introduce the NAH field theories (NAHFTs), i.e. field theories describing multiflavor scalar fields coupled with non-Abelian gauge fields. In particular, we discuss statistical field theories with local SU($N_{c}$) and SO($N_{c}$) gauge symmetry, and scalar fields associated with different representations of the gauge group, such as the fundamental and adjoint representation. We outline the known features of their RG flow, whose stable infrared FPs are expected to describe the charged critical behaviors of the corresponding statistical systems.
@@ 8.1.1 The statistical field theory
We consider the NAHFT in which SU($N_{c}$) gauge fields interact with $N_{f}$ degenerate scalar fields in the fundamental representation of the gauge group. The fundamental fields are a complex scalar field $\Phi^{af}(\bm{x})$, where $a=1,...,N_{c}$ and $f=1,...,N_{f}$, and an SU($N_{c}$) gauge field ${\cal A}_{\mu}=\sum_{c=1}^{N_{c}^{2}-1}A_{\mu}^{c}T^{c}$, where $T^{c}$ are the Hermitian generators of the fundamental representation of SU($N_{c}$), normalized as $\mathrm{Tr}(T^{a}T^{b})=\frac{1}{2}\delta^{ab}$. The most general renormalizable Lagrangian that is invariant under local SU($N_{c}$) color transformations and under global U($N_{f}$) flavor transformations reads [273, 99]
@@
where ${\cal F}_{\mu\nu}=\partial_{\mu}{\cal A}_{\nu}-\partial_{\nu}{\cal A}_{\mu}-i[{\cal A}_{\mu},{\cal A}_{\nu}]$, and $D_{\mu,ab}$ is the covariant derivative associated with the fundamental representation, i.e., $D_{\mu,ab}=\partial_{\mu}\delta_{ab}-i\sum_{c}T_{ab}^{c}A_{\mu}^{c}$. The global U($N_{f}$) group acts as $\Phi({\bm{x}})\to\Phi(\bm{x})W$, with $W\in\mathrm{U}(N_{f})$, while the local SU($N_{c}$) gauge group acts as
@@
We finally note that, for $N_{c}=2$, because of the pseudoreality of the SU(2) gauge group, the NAHFT (85) with $v=0$ turns out to be invariant under transformations associated with a symmetry group that is larger than U($N_{f}$), i.e. the Sp($N_{f}$)/${\mathbb{Z}}_{2}$ group [370, 379, 375, 32, 97, 167, 99]. In particular, it is invariant under Sp(2)/${\mathbb{Z}}_{2}$ = SO(5) transformations for $N_{f}=2$ [97, 167]. Therefore, the $N_{c}=2$ NAHFT (85) with $v=0$ is stable under RG transformations, defining a distinct theory characterized by the global symmetry Sp($N_{f}$)/${\mathbb{Z}}_{2}$.
@@ 8.1.2 Higgs phases and breaking patterns of the global symmetry
The spontaneous global symmetry breaking of the Higgs phases, which is nontrivial only for $N_{f}\geq 2$, can be studied by performing an analysis of the minima of the scalar potential within a semiclassical approximation. We report here the main results. See Ref. [99] for more details, and C for a sketch of the derivations in the corresponding lattice models. The results of such mean-field analyses depend on the sign of $v$ and, for $v>0$, also on the number of colors and flavors, i.e., $N_{c}$ and $N_{f}$.
@@
For $v>0$ one should distinguish three different cases. For $N_{f}<N_{c}$ there is no symmetry breaking and therefore, no Higgs ordered phase. For $N_{f}\geq N_{c}$ we have instead
@@ 8.1.3 Renormalization-group flow
A nonperturbative formulation of the NAHFT can be obtained as the continuum (critical) limit of corresponding lattice models with analogous local gauge and global symmetries, as it is done for the strong-interaction theory, the quantum chromodynamics (QCD) [100, 3]. For this purpose it is crucial that the corresponding lattice discretizations has a continuous charged transition where the gauge fields play a relevant role. Therefore, it is important to identify the universality classes of the charged transitions in this class of systems. They must be related to the RG flow of the NAHFT, and in particular to its stable FPs.
The RG flow of the NAHFT (85) can be investigated perturbatively, in powers of $\varepsilon\equiv 4-d$ [94], using dimensional regularization and the minimal-subtraction (MS) renormalization scheme, see, e.g., Ref. [2, 95]. The RG flow close to four dimensions is determined by the MS $\beta$ functions associated with the Lagrangian couplings $\alpha=g^{2}$, $u$, and $v$. One-loop $O(\varepsilon)$ computations give [99, 273]
@@
Some numerical factors have been reabsorbed in the normalizations of the renormalized couplings, in order to simplify the expressions of their $\beta$ functions.
The analysis of the common zeroes of the $\beta$ functions shows that the RG flow close to four dimensions has a stable charged FP with a nonzero FP value of the gauge coupling $\alpha$, for a sufficiently large number of flavors; more precisely, for $N_{f}>N_{f}^{*}$, with, e.g., $N_{f}^{*}=375.4+O(\varepsilon)$ for $N_{c}=2$, and $N_{f}^{*}=638.9+O(\varepsilon)$ for $N_{c}=3$. The values of $N_{f}^{*}$ are large in four dimensions, but they are expected to decrease significantly in three dimensions for any $N_{c}$, as it also happens in the AHFT, where $N_{f}^{*}$ varies from $N_{f}^{*}\approx 183$ in four dimensions to $N_{f}^{*}\approx 7$ in three dimensions. In particular, for $N_{c}=2$ numerical studies based on MC simulations of lattice models [392] indicate $N_{f}^{*}=25(4)$ in three dimensions, see Sec. 9.3. We also mention that in the case of the SU(2)-gauge NAHFT (85) for $v=0$, for which the global symmetry enlarges to Sp($N_{f}$)/${\mathbb{Z}}_{2}$, a charged FP exists for $N_{f}>N_{f}^{*}\approx 359+O(\varepsilon)$ [97]. This FP turns out to be unstable in the full theory, i.e, with respect to nonzero values of the Lagrangian coupling $v$.
The stable charged FP of this class of NAHFTs is located in the positive $v$ region for any $N_{c}$. This result indicates that only transitions characterized by the $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$ symmetry-breaking pattern may correspond to charged continuous transitions. Those characterized by the symmetry-breaking pattern $\mathrm{U}(N_{f})\to{\mathrm{U}}(1)\oplus{\mathrm{U}}(N_{f}-1)$, corresponding to $v<0$, instead cannot be charged continuous transitions. As we have already discussed for the AH models, it is crucial to note that the FPs of the NAHFTs are relevant for the existence of charged continuous transitions, i.e., of continuous transitions with critical behaviors where both scalar and gauge fields play a role. As discussed in Sec. 3.1, gauge models can also undergo other types of transitions that admit different effective descriptions. Therefore, it is a priori possible to have further continuous transitions, which may be LGW or topological transitions, also for $N_{f}<N_{f}^{*}$ and for $v<0$. Some examples will be discussed in Sec. 9.
As it occurs in the models with Abelian gauge symmetry, the uncharged FPs with vanishing gauge coupling $\alpha=0$ are unstable with respect to the gauge coupling, since their stability matrix $\Omega_{ij}=\partial\beta_{i}/\partial g_{j}$ has a negative eigenvalue [99]
@@
This result suggests that also in three dimensions the gauge fluctuations makes the uncharged FP unstable, as also confirmed by the numerical studies of 3D SU(2)-gauge LNAH models [99], see also Sec. 9. Note that continuous transitions described by the uncharged FP of the NAHFT for $\alpha=0$ would be characterized by an asymptotic critical behavior in which gauge-field correlations are suppressed, thus they would correspond to LGW× transitions, effectively described by a LGW theory with a gauge-dependent complex order-parameter field $\Phi^{af}(\bm{x})$ with $a=1,...,N_{c}$ and $f=1,...,N_{f}$.

### v3

@@ 7.5.2 Gauge symmetry enlargement
As we have already discussed several times, see, e.g., Sec. 7.5.1, at transitions of LGW and LGW× type, gauge models may display an enlargement of the global symmetry. Analogously, at topological transitions, one may observe an enlargement of the gauge symmetry—this occurs in ${\mathbb{Z}}_{Q}$ gauge models, see Sec. 3.5.2. The models we are now going to discuss have instead been investigated to understand whether there are GFT transitions, effectively described by a continuum GFT, that show a gauge-symmetry enlargement. In the present context, one is looking for models with discrete gauge symmetry that undergo charged transitions that are associated with the AHFT, with an enlarged U(1) gauge symmetry.
Charged transitions controlled by the AHFT occur in the $N$-component LAH model with noncompact Abelian gauge field (along the CH line) and in the higher-charge ($Q>1$) $N$-component LAH model with compact Abelian gauge field (along the DC-OD line). In both cases the gauge group can be approximated with arbitrary accuracy by considering discrete compact or noncompact subgroups and one can define models that are only invariant under these discrete subgroups. The question is whether these discrete models also have charged transitions as their continuous counterparts.
We begin by discussing the discretized version [401] of the noncompact $N$-component LAH model with Hamiltonian (47) defined in Sec. 5.1. The discretized model is defined in the same way, the only difference being the set of values that the gauge field $A_{{\bm{x}},\mu}$ takes. In the discretized model, the gauge field takes only the values $2\pi m/q$, with $m\in\mathbb{Z}$. The model is invariant under a restricted set of gauge transformations (we indicate with ${\mathbb{Z}}_{q}^{(nc)}$ the corresponding group)
@@
The original noncompact U(1) gauge theory is formally recovered in the limit $q\to\infty$. Note that the gauge group is discrete, but infinite. As already discussed in Sec. 5.1, a proper definition of the model requires the introduction of $C^{*}$ boundary conditions.
Figure 18: Sketch of the phase diagram of the 3D $N$-component LAH model with discrete gauge symmetry. In the noncompact case the transition in $A$ belongs to the XY universality class, while the transitions at points $B$ and $C$ belong to the IXY universality class [401]. In the compact case the transition in $A$ is in the XY universality class, while the transitions at points $B$ and $C$ are in the $\mathbb{Z}_{Q}$ and $\mathbb{Z}_{q}$ universality class, respectively [241]. Both in the compact and in the noncompact case, we have $J_{c}(A),\kappa_{c}(C)\to\infty$ in the limit $q\to\infty$.
@@
Due to its larger global symmetry, the phase diagram of the discrete $N$-component LAH model is significantly more complex than that of the $N$-component LAH with U(1) gauge symmetry. A sketch is shown in Fig. 18. Using the same arguments presented in Sec. 6.2, it is possible to show (see Ref. [401] for more details) that an IXY transition occurs both for $J=\infty$ (as in the noncompact U(1) gauge AH model) and for $J=0$. Moreover, the transition on the $J=0$ axis occurs at $\kappa_{c}(J=0)=q^{2}\kappa_{c}(J=\infty)$, where $\kappa_{c}(J=\infty)=0.076051(2)$, so $\kappa_{c}(J=0)\to\infty$ for $q\to\infty$. This should not come as a surprise, since we should recover the phase diagram of the model with U(1) symmetry with no transition on the $J=0$ line, see Fig. 2, for $q\to\infty$. Moreover, two transitions are present on the $\kappa=0$ axis: a transition for $J=J_{\mathrm{SU}(N)}$, where the global SU($N$) symmetry is spontaneously broken, and a transition at $J=J_{q}$, associated with the spontaneous breaking of the U(1)$/\mathbb{Z}_{q}$ symmetry. As discussed above, we expect the latter symmetry to break in a phase in which the SU($N$) symmetry is already broken, implying $J_{q}>J_{\mathrm{SU}(N)}$ (therefore the point $A$ in Fig. 18 corresponds to $J=J_{q}$). Note that $J_{q}\to\infty$ as $q\to\infty$, since this transition is not present in the phase diagram of the noncompact U(1) gauge model, see Fig. 2.
In Ref. [401] the phase diagram of the $N$-component $\mathbb{Z}_{q}^{(nc)}$ gauge model has been numerically investigated, studying in detail the critical behavior along the transition line $M_{2}M_{3}$ in Fig. 18. Indeed, for $q\to\infty$, this line corresponds to the CH line (see Fig. 2) of the noncompact U(1) gauge model, which is the line where the AHFT critical behavior emerges if the number $N$ of scalar components is large enough, as discussed in Sec. 5.3.3. The numerical results show that, for $N=25$ and $q\geq 5$, the transitions along the line $M_{2}M_{3}$ are continuous, and that their critical behavior is the same as along the CH line in the noncompact U(1) gauge model. Thus, for $q\geq 5$ and large values of $N$, the transitions along the line $M_{2}M_{3}$ are effectively associated with the stable charged FP of the AHFT, with a corresponding enlargement of the gauge symmetry. On the right of the intersection point $M_{3}$ we instead have O($2N$)× critical transitions.
@@
In Ref. [241] the universality class of the transitions along the line $M_{2}M_{3}$ has been investigated for $Q=2$ and $N=15$. In the U(1) invariant model, recovered for $q\to\infty$, these transitions are charged transitions associated with the stable FP of the AHFT, see Sec. 6.3.2. On the other hand, only discontinuous transitions were identified along the $M_{2}M_{3}$ line for values of $q$ up to $q=10$. Apparently, in the discretized model there are no charged transitions. It is still an open issue to understand why the two apparently similar discretized models, the compact and the noncompact model, behave differently. The origin may lie in the fact that in the noncompact model the gauge group is discrete but still infinite, while in the compact case the gauge group is discrete and finite.
@@ 8 Non-Abelian Higgs field theories
Non-Abelian gauge symmetries play a fundamental role in the construction of quantum and statistical field theories that describe phenomena in various physical contexts. In high-energy physics they provide the framework to formulate the fundamental theories of the strong and electroweak interactions, see, e.g., Refs. [1, 2, 402], while in condensed-matter physics they provide an effective description of classical and quantum critical phenomena characterized by the emergence of non-Abelian gauge fields, see, e.g., Refs [403, 97, 20, 404, 16]. In the presence of low-energy modes associated with scalar fields, they give rise to the non-Abelian Higgs mechanism [104, 105, 106], which, for instance, is one of the main features of the theory of electroweak interactions [1, 2, 402]. The emerging Higgs phases in non-Abelian Higgs (NAH) theories crucially depend on the interplay between local and global symmetries, and on the pattern of the spontaneous symmetry breaking, see, e.g., Refs. [405, 288, 41, 406, 304, 62, 64, 65, 33, 407, 97, 98, 289, 408, 99, 100].
Three-dimensional non-Abelian gauge models can be used to effectively describe the critical behavior of 3D statistical non-Abelian gauge models at finite-temperature phase transitions [409, 410, 411, 412, 413, 414, 415, 416, 417, 418], and to investigate some aspects of fundamental mechanisms, such as confinement [405, 419, 420, 286]. The peculiar critical behaviors occurring in 3D NAH models also provide additional possible scenarios for finite-temperature continuous transitions in $D=3+1$ non-Abelian gauge QFTs. In this case, temperature is related to the (inverse) finite size along the fourth Euclidean direction in the corresponding path-integral formulation, see, e.g., Refs. [25, 409, 410, 411, 412, 413, 414, 421, 415, 253, 254, 199, 26, 23, 200]. NAH theories have also been considered to explain some aspects of the rich phenomenology of high-temperature cuprate superconductors near optimal doping, with display the simultaneous disappearance of nematic and charge-density orders, the reconstruction of the Fermi surface, and the restoration of time-reversal symmetry; see, e.g., Refs. [422, 423, 424, 425], and Ref. [426] for a review. In particular, the possible emergence at criticality of a non-Abelian gauge symmetry in the presence of scalar fields has been proposed in Refs. [403, 97]; see also Refs. [427, 190, 428, 429, 430, 431, 432] for earlier studies.
In this section we introduce the NAH field theories (NAHFTs), i.e., field theories in which scalar fields are coupled with non-Abelian gauge fields. In particular, we discuss statistical field theories with local SU($N_{c}$) and SO($N_{c}$) gauge symmetry, and scalar fields transforming according to different representations of the gauge group (we consider the fundamental and the adjoint representation). We outline the known features of the RG flow, whose stable infrared FPs are expected to describe the charged critical behaviors occurring in the corresponding lattice systems.
@@ 8.1.1 The statistical field theory
We consider the NAHFT in which SU($N_{c}$) gauge fields interact with $N_{f}$ degenerate scalar fields in the fundamental representation of the gauge group. The fundamental fields are a complex scalar field $\Phi^{af}(\bm{x})$, where $a=1,...,N_{c}$ and $f=1,...,N_{f}$, and the gauge field $A_{\mu}^{c}({\bm{x}})$, where $c=1,...,N_{c}^{2}-1$. The most general renormalizable Lagrangian that is invariant under local SU($N_{c}$) color transformations and under global U($N_{f}$) flavor transformations reads [304, 100]
@@
where ${\cal F}_{\mu\nu}=\partial_{\mu}{\cal A}_{\nu}-\partial_{\nu}{\cal A}_{\mu}-i[{\cal A}_{\mu},{\cal A}_{\nu}]$, ${\cal A}_{\mu}=\sum_{c}A_{\mu}^{c}T^{c}$, $D_{\mu}$ is the covariant derivative in the fundamental representation, i.e., $D_{\mu,ab}=\partial_{\mu}\delta_{ab}-i\sum_{c}T_{F,ab}^{c}A_{\mu}^{c}$, where $T_{F}^{c}$ are the $N_{c}^{2}-1$ Hermitean generators in the fundamental representation of the SU($N_{c}$) algebra, normalized as $\mathrm{Tr}(T^{a}T^{b})=\frac{1}{2}\delta^{ab}$. The global U($N_{f}$) group acts as $\Phi({\bm{x}})\to\Phi(\bm{x})W$, with $W\in\mathrm{U}(N_{f})$, while the local SU($N_{c}$) gauge group acts as
@@
For $N_{c}=2$, because of the pseudoreality of the SU(2) gauge group, the NAHFT Hamiltonian (88) with $v=0$ turns out to be invariant under Sp($N_{f}$)/${\mathbb{Z}}_{2}$ transformations, i.e., under a group that is larger than U($N_{f}$), [402, 411, 407, 32, 98, 198, 100]. In particular, it is invariant under Sp(2)/${\mathbb{Z}}_{2}$ = SO(5) transformations for $N_{f}=2$ [98, 198]. Therefore, the $N_{c}=2$ NAHFT (88) with $v=0$ is stable under RG transformations, defining a distinct theory characterized by the global symmetry Sp($N_{f}$)/${\mathbb{Z}}_{2}$.
We finally recall that the standard Faddeev-Popov quantization [433] of a non-Abelian gauge field theory requires the addition of a gauge fixing. If we consider the Lorenz gauge fixing used in the AHFT defined in Eq. (22), one should also introduce appropriate scalar fields with fermionic statistics (ghosts), see, e.g., Refs. [1, 2].
@@ 8.1.2 Higgs phases and breaking patterns of the global symmetry
The global symmetry of the Higgs phases, which is nontrivial only for $N_{f}\geq 2$, can be studied by performing an analysis of the minima of the scalar potential. We report here the main results. See Ref. [100] for more details, and C for a sketch of the derivations in the corresponding lattice models. The results of these analyses depend on the sign of $v$ and, for $v>0$, also on the number of colors $N_{c}$ and flavors $N_{f}$.
@@
For $v>0$ one should distinguish three different cases. For $N_{f}<N_{c}$ there is no symmetry breaking and therefore no Higgs ordered phase. For $N_{f}\geq N_{c}$ we have instead
@@ 8.1.3 Renormalization-group flow
A nonperturbative formulation of the NAHFT is provided by the continuum (critical) limit of corresponding lattice models with the same local and global symmetries, as it is done for the strong-interaction theory, quantum chromodynamics (QCD) [101, 3]. For this purpose it is crucial that the corresponding lattice discretizations have a continuous charged transition where the gauge fields play a relevant role. Therefore, it is important to identify the universality classes of the charged transitions in this class of systems. They are related to the stable FPs of the RG flow of the NAHFT.
The RG flow of the NAHFT (88) can be investigated perturbatively, in powers of $\varepsilon\equiv 4-d$ [95], using dimensional regularization and the minimal-subtraction (MS) renormalization scheme, see, e.g., Ref. [2, 96]. The RG flow close to four dimensions is determined by the MS $\beta$ functions associated with the Lagrangian couplings $\alpha=g^{2}$, $u$, and $v$. At one-loop order, we have [100, 304]
@@
Some numerical factors have been reabsorbed in the normalizations of the renormalized couplings, in order to simplify the previous expressions.
The analysis of the common zeroes of the $\beta$ functions shows that the RG flow close to four dimensions has a stable charged FP with a nonzero FP value of the gauge coupling $\alpha$, for a sufficiently large number of flavors; more precisely, for $N_{f}>N_{f}^{*}$, with, e.g., $N_{f}^{*}=375.4+O(\varepsilon)$ for $N_{c}=2$, and $N_{f}^{*}=638.9+O(\varepsilon)$ for $N_{c}=3$. The values of $N_{f}^{*}$ are large in four dimensions, but they are expected to decrease significantly in three dimensions for any $N_{c}$, as it also happens in the AHFT, where $N_{f}^{*}$ varies from $N_{f}^{*}\approx 183$ in four dimensions to $N_{f}^{*}\approx 7$ in three dimensions. In particular, for $N_{c}=2$ numerical MC studies of lattice SU(2)-gauge Higgs models [434] indicate $N_{f}^{*}=25(4)$ in three dimensions, see Sec. 9.3. We also mention that in the case of the SU(2)-gauge NAHFT (88) for $v=0$, for which the global symmetry enlarges to Sp($N_{f}$)/${\mathbb{Z}}_{2}$, a charged FP exists for $N_{f}>N_{f}^{*}\approx 359+O(\varepsilon)$ [98]. This FP turns out to be unstable in the full theory, i.e, with respect to nonzero values of the Lagrangian coupling $v$.
The stable charged FP of this class of NAHFTs is located in the positive $v$ region for any $N_{c}$. This result indicates that only transitions characterized by the $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$ symmetry-breaking pattern may correspond to charged continuous transitions. Instead, those characterized by the symmetry-breaking pattern $\mathrm{U}(N_{f})\to{\mathrm{U}}(1)\oplus{\mathrm{U}}(N_{f}-1)$, corresponding to $v<0$, cannot be charged continuous transitions. As we have already discussed for the AH models, the FPs of the NAHFTs are relevant for the existence of charged continuous transitions, i.e., of continuous transitions where both scalar and gauge fields are critical. As discussed in Sec. 3.1, gauge models can also undergo other types of transitions that admit different effective descriptions. Therefore, it is a priori possible to have continuous transitions, which may be LGW or topological transitions, also for $N_{f}<N_{f}^{*}$ and for $v<0$. Some examples will be discussed in Sec. 9.
As it occurs in models with Abelian gauge symmetry, the uncharged FP with vanishing gauge coupling $\alpha=0$ is unstable with respect to the gauge coupling, since its stability matrix $\Omega_{ij}=\partial\beta_{i}/\partial g_{j}$ has a negative eigenvalue [100]
@@
This result suggests that also in three dimensions gauge fluctuations make the uncharged FP unstable, as also confirmed by the numerical studies of 3D SU(2)-gauge LNAH models [100], see also Sec. 9. Note that continuous transitions described by the uncharged FP of the NAHFT for $\alpha=0$ would be characterized by an asymptotic critical behavior in which gauge-field correlations are not critical, thus they would correspond to LGW× transitions, effectively described by a LGW theory with a gauge-dependent complex order-parameter field $\Phi^{af}(\bm{x})$ with $a=1,...,N_{c}$ and $f=1,...,N_{f}$.